### v1

# UniSite: The First Cross-Structure Dataset and Learning Framework for End-to-End Ligand Binding Site Detection

Abstract
The detection of ligand binding sites for proteins is a fundamental step in Structure-Based Drug Design. Despite notable advances in recent years, existing methods, datasets, and evaluation metrics are confronted with several key challenges: (1) current datasets and methods are centered on individual protein–ligand complexes and neglect that diverse binding sites may exist across multiple complexes of the same protein, introducing significant statistical bias; (2) ligand binding site detection is typically modeled as a discontinuous workflow, employing binary segmentation and subsequent clustering algorithms; (3) traditional evaluation metrics do not adequately reflect the actual performance of different binding site prediction methods. To address these issues, we first introduce UniSite-DS, the first UniProt (Unique Protein)-centric ligand binding site dataset, which contains 4.81 times more multi-site data and 2.08 times more overall data compared to the previously most widely used datasets. We then propose UniSite, the first end-to-end ligand binding site detection framework supervised by set prediction loss with bijective matching. In addition, we introduce Average Precision based on Intersection over Union (IoU) as a more accurate evaluation metric for ligand binding site prediction. Extensive experiments on UniSite-DS and several representative benchmark datasets demonstrate that IoU-based Average Precision provides a more accurate reflection of prediction quality, and that UniSite outperforms current state-of-the-art methods in ligand binding site detection. The dataset and codes will be made publicly available at https://github.com/quanlin-wu/unisite.
11footnotetext: Equal contribution.22footnotetext: Corresponding author.
1 Introduction
The detection of ligand binding sites on target proteins is one of the most critical steps in modern drug discovery strategies nakashima2011structures ; vincent2022phenotypic ; chan2019new . Structure-based drug design approaches begin with the three-dimensional structure of the target protein, from which deep, druggable cavities are identified. These regions, referred to as binding sites or binding pockets, are composed of sets of protein residues. Once the protein’s sites are recognized, virtual screening of a molecular library can be performed using methods such as protein–ligand docking and protein–ligand affinity prediction campbell2003ligand ; li2023pocketanchor . Alternatively, de novo molecular design powers2023geometric ; schneuing2024structure can be conducted based on the local structure of the binding sites to identify potential candidate compounds. As a fundamental step, the accurate identification of protein binding sites can significantly facilitate and influence subsequent steps in drug discovery.
Over the past several decades, some endeavours have been made to detect protein ligand binding sites. These methods have evolved from traditional techniques based on geometry fpocket2009 , template searching findsite2008 , and energy probes ngan2012ftsite , to machine learning methods based on surface features p2rank2018 , and further to deep learning methods utilizing convolutional neural networks (CNNs) deeppocket2022 and graph neural networks (GNNs) equipocket2023 ; grasp2024 ; vnegnn2024 . Concurrently, a series of protein–ligand datasets have also been established progressively, including scPDB scpdb2015 and PDBbind pdbbind2004 datasets for protein–ligand complex structures, as well as benchmark datasets such as HOLO4K schmidtke2010large and COACH420 yang2013protein for evaluating binding site detection methods.
Although the above efforts have significantly advanced the field of ligand binding site detection, current methods, datasets, and evaluation metrics are confronted with substantial challenges:
Issue 1. All previous methods and datasets are PDB (Protein Data Bank file)-centric, specifically focusing on individual protein–ligand structures, which introduces considerable statistical bias. Due to experimental constraints, only a limited number of binding sites in the protein are typically observed in one single protein–ligand structure where ligands are bound. However, one protein can be associated with numerous distinct protein–ligand structures, which exhibit high structural similarity in their protein components yet considerable variation in their binding site regions he2024allosteric ; matthees2024ca2 ; morstein2024targeting (Figure 1). But existing datasets and methods only regard these structures as individual data entries, focusing on limited binding sites in single PDB structure. Training and evaluating on PDB-centric datasets introduces significant statistical bias, as the annotation paradigm of individual PDB structures overlooks many other ground truth binding sites.
Issue 2. Existing methods employ discontinuous workflows for binding site detection. Most approaches fpocket2009 ; p2rank2018 ; equipocket2023 ; grasp2024 first perform semantic segmentation to generate binary masks of potential binding residues/atoms, then cluster them into discrete binding sites. Alternative implementations only predict binding site centers vnegnn2024 , and the associated residues need to be extracted using external methods. These fragmented pipelines highly rely on the post-processing methods (e.g. clustering algorithms), inherently limit end-to-end optimization and struggle with overlapping binding sites.
Issue 3. Traditional evaluation metrics inadequately reflect the actual performance of binding site detection. The most widely used evaluation metrics are DCC and DCA p2rank2018 . DCC represents the distance between the predicted binding site center and the ground truth binding site center. DCA denotes the shortest distance between the predicted binding site center and any heavy atom of the ligand. These metrics suffer from two fundamental limitations (Figure 4): (1) they completely disregard the structural properties such as shape, size, and residue composition of binding sites, which are crucial for downstream tasks (Appendix A), and (2) the absence of proper matching criteria between predictions and ground truth may lead to double-counting of predictions.
To address the issues mentioned above, this paper makes the following contributions:
1) We introduce UniSite-DS, a manually curated, UniProt (Unique Protein)-centric dataset of protein ligand binding sites. Leveraging the unique identifiers assigned to protein sequences in UniProt uniprot2019uniprot , we systematically integrated all ligand binding sites associated with given unique protein across multiple PDB structures. To the best of our knowledge, it is the first UniProt-centric dataset. Notably, UniSite-DS includes 4.81 times more multi-site proteins than existing datasets scpdb2015 ; pdbbind2004 , and the overall size of the dataset is 2.08 times larger. The Uniprot-centric dataset corrects the statistical bias of previous PDB-centric datasets, thereby resolving Issue 1 and significantly broadening the available data.
2) We propose UniSite-1D and UniSite-3D, two end-to-end methods for protein ligand binding site detection. Both models utilize a transformer encoder-decoder architecture, supervised by a set prediction loss with bijective matching. UniSite 1D/3D directly predict $N$ potentially overlapping binding sites without requiring post-processing clustering steps, thus completely resolves Issue 2. The UniSite-1D variant operates exclusively on 1D protein sequence inputs, providing structure-free binding site detection capability. For enhanced performance, the UniSite-3D variant incorporates 3D structural information while maintaining the same end-to-end prediction framework.
3) To overcome the limitations inherent in traditional evaluation methods outlined in Issue 3, we introduce an Average Precision (AP) metric based on Intersection over Union (IoU) for comprehensive binding site assessment. Extensive experiments have demonstrated that the IoU-based AP maintains strong concordance with method rankings under traditional metrics while overcoming their key limitations, providing a more accurate reflection of prediction quality.
4) Extensive experiments on UniSite-DS and classical datasets have demonstrated that our methods outperform the current state-of-the-art methods in protein ligand binding site detection. These results indicate that the end-to-end detection framework, which operates without the need for specialized feature engineering, is already capable of exhibiting strong performance for binding site detection.
Figure 1: Comparison between UniSite-DS and previous datasets. (Top left) In PDBbind2020, only one ligand binding site and one structure are recorded for UniProt ID Q8WS26. (Top right) In contrast, UniSite-DS integrates distinct binding sites across all available structures (highly similar, mean TM-Score=0.99), identifying 17 unique ligand binding sites derived from 13 representative PDB entries. (Bottom left and center) Comparison of UniSite-DS with other widely used datasets in terms of multi-site entries and the number of unique proteins. For HOLO4K and COACH420, the most widely used mlig subsets were selected, where each entry corresponds to a PDB structure, while in UniSite-DS, each entry corresponds to a UniProt ID. (Bottom right) Distribution of the number of unique proteins in UniSite-DS with respect to the number of distinct binding sites they contain.
2 UniSite-DS: The First Uniprot-centric Dataset
A key challenge in detecting protein binding sites is how to identify all potential binding sites he2024allosteric ; matthees2024ca2 . Most proteins contain an inherently conserved binding site, commonly referred to as the active site. The active site is shared among members of the same protein family, which means that molecules targeting this site will simultaneously target all other proteins within the family, which is highly likely to lead to off-target effects and side effects zhang2024recent . Identifying other binding sites within the protein that can be targeted is a crucial strategy. These sites are often located in regions topologically distant from the active site and can modulate the protein’s function through allosteric effects hughes2019asciminib ; slosky2020beta ; fan2021harnessing .
As illustrated in Figure 1, one single protein can correspond to a large number of different ligand-bound structures. While the overall protein structure tends to be highly conserved, the ligand binding site regions vary considerably across these structures. The motivation behind constructing UniSite-DS lies in the recognition that identifying all potential binding sites of a protein requires a comprehensive examination of all its ligand-bound structures—an important consideration that has been overlooked by previous methods and datasets.
To construct UniSite-DS, we performed the following search and processing steps: (1) We utilized AHoJ feidakis2022ahoj to systematically search for all protein–ligand interactions in the PDB database berman2003announcing ; (2) To ensure dataset quality, we excluded entries with a resolution greater than 2.5Å or those determined by non-crystallographic methods; (3) Following P2Rank’s filtering criteria p2rank2018 , we removed solvent molecules and ligands composed of fewer than five atoms, resulting in a total of 143,197 protein–ligand interaction entries; (4) For each interaction, binding site residues were identified within a 4.5Å radius of the ligand; (5) We discarded entries with three or fewer binding site residues to eliminate “floating” ligands; (6) Leveraging UniProt’s unique protein sequence identifiers uniprot2019uniprot , we mapped binding site residues from all protein–ligand interactions of each UniProt entry to their corresponding sequences via SIFTS annotations velankar2012sifts , integrating all ligand binding sites across different PDB structures; (7) To eliminate data redundancy among ligand binding sites, we applied Non-Maximum Suppression (NMS) with an Intersection over Minimum (IoM) threshold of 0.7 and an Intersection over Union (IoU) threshold of 0.5, excluding highly overlapping sites. This process resulted in 13,464 distinct UniProt IDs, of which 4,846 contained multiple ligand-binding sites; (8) Based on the criteria from Proteina geffner2025proteina , we set the sequence length threshold to 800; (9) We manually inspected all UniProt IDs with more than ten ligand binding sites, as well as those where a single protein–ligand complex structure contributed three or more binding sites. As a result, we identified 11,510 valid UniProt IDs, including 3,670 with multiple ligand binding sites. The distribution of ligand binding sites is shown in Figure 1.
As the first UniProt-centric dataset, UniSite-DS encompasses 4.81 times more multi-site entries than previous datasets, and covers 2.96 times more UniProt entries than the widely used PDBbind dataset pdbbind2004 , as well as 2.08 times more than sc-PDB scpdb2015 (Figure 1). UniSite-DS eliminates the statistical biases inherent in earlier datasets and significantly expands the available data on multi-site ligand binding sites. Notably, case studies conducted using UniSite-DS (Appendix B) highlighted the limitations of current binding site prediction methods in handling multi-site proteins. This observation motivated us to develop a novel end-to-end method for protein–ligand binding site detection.
3 The Proposed Methodology
Figure 2: Comparison of detection approaches. (Top) Conventional learning-based binding site detection methods typically employ a discontinuous workflow: first predicting binary masks for residues/atoms, then clustering these masks into distinct binding sites. (Bottom) In contrast, our method directly outputs a set of $N$ potentially overlapping binding sites in a single step.
In this paper, we formulate protein ligand binding site detection as a set prediction task: given a protein $P$ with an amino acid sequence $S$ of length $L$, the goal of binding site detection is to identify a set of binding sites $\{m_{i}^{gt}\}_{i=1}^{N_{gt}}$, where each binding site is represented by a binary mask $m_{i}^{gt}\in\{0,1\}^{L}$. Here, $m_{ij}^{gt}=1$ indicates that the $j$-th residue is part of the $i$-th site, while $m_{ij}^{gt}=0$ means it is not. Currently, most learning-based binding site detection methods adopt a discontinuous workflow: first predicting a score for each amino acid residue or heavy atom, and then clustering them into distinct binding sites. To streamline this process, we propose UniSite, the first UniProt-centric and direct set prediction approach that adheres to the end-to-end paradigm (Figure 2). Two components are essential for direct set prediction in this context: (1) a set prediction loss based on bijective matching between predicted and ground truth binding sites; and (2) an architecture capable of predicting a set of sites in a single forward pass. The architecture of UniSite is shown in detail in Figure 3.
3.1 Set prediction loss for binding site detection
UniSite infers a fixed-size set of $N$ predictions $z=\{(p_{i},m_{i})|m_{i}\in\{0,1\}^{L}\}_{i=1}^{N}$ in a single forward pass, where $m_{i}$ represents the predicted binding site, and $p_{i}$ denotes the probability of binding and $\emptyset$ (non-binding) category. Since the ground truth set $|z^{gt}|=N^{gt}$ and the prediction set $|z|=N$ typically have unequal sizes, we assume $N\geq N^{gt}$ and pad the ground truth set with $\emptyset$ (non-binding) tokens. The padded ground truth set is defined as $z^{gt}_{pad}=\{(c_{i}^{gt},m_{i}^{gt})|c_{i}^{gt}\in\{1,\emptyset\},m_{i}^{gt}\in\{0,1\}^{L}\}_{i=1}^{N}$, where $c_{i}^{gt}=1$ indicates a true binding site and $c_{i}^{gt}=\emptyset$ corresponds to the padding.
To train a set prediction model, we require a bijective matching $\sigma$ between the predicted set $z$ and the padded ground truth set $z^{gt}_{pad}$. This bipartite matching is obtained by minimizing the following cost:
[…]
Figure 3: The architecture of UniSite. Our models employ an encoder to extract the residue-level features. Then a decoder module is used to generate embeddings of the $N$ predicted binding sites. Finally, the segmentation module outputs $N$ potentially overlapping binding site predictions. The encoder comprises dual pathways: a sequence encoder and an optional structural encoder, allowing UniSite to operate with either sequence-only input or combined sequence-structure information.
where $\sigma$ is a permutation of $N$ elements and $\mathcal{L}_{\text{match}}$ quantifies the pairwise matching cost between ground truth site $z_{i}^{gt}$ and the prediction with index $\sigma(i)$. Following prior work (detr2020, ; Stewart2016, ), we employ the Hungarian algorithm to compute the optimal matching and use the matching cost defined as:
$\mathcal{L}_{\text{match}}(z_{i}^{gt},z_{\sigma(i)})=-\mathbf{1}_{\{c_{i}^{gt}\neq\emptyset\}}\text{log}p_{\sigma(i)}(c_{i}^{gt})+\mathbf{1}_{\{c_{i}^{gt}\neq\emptyset\}}\mathcal{L}_{\text{mask}}(m_{i}^{gt},m_{\sigma(i)})$
where $\mathcal{L}_{\text{mask}}$ is a binary mask loss, $p_{\sigma(i)}$ and $m_{\sigma(i)}$ denote the predicted probability and the binding site for the $\sigma(i)$-th prediction, respectively. This matching cost considers both the class prediction and the similarity of the predicted and ground truth binding sites. Given the optimal matching $\hat{\sigma}$, we compose a cross-entropy classification loss and a binary mask loss $\mathcal{L}_{\text{mask}}$ for each predicted site to train model parameters:
$\mathcal{L}_{\text{mask\&cls}}(z^{gt},z)=\sum_{i}^{N}{-\log p_{\hat{\sigma}(i)}(c_{i}^{gt})}+\mathbf{1}_{\{c_{i}^{gt}\neq\emptyset\}}\mathcal{L}_{\text{mask}}(m_{i}^{gt},m_{\hat{\sigma}(i)})$
3.2 UniSite architecture
As illustrated in Figure 3, UniSite comprises three main components: (1) an encoder module that extracts residue-level representations $\mathcal{F}\in\mathbb{R}^{L\times d_{\text{model}}}$ of the protein; (2) a decoder module consisting of multiple Transformer decoder layers to generate embeddings of the $N$ predicted binding sites, and (3) a segmentation module which combines the residue-level representations and the decoder embeddings to produce the final predictions $\{(p_{i},m_{i})|m_{i}\in\{0,1\}^{L}\}_{i=1}^{N}$. This architecture maintains conciseness while demonstrating strong compatibility with existing protein representation methods. In our implementation, we construct two variants: UniSite-1D using only sequence encoding, and UniSite-3D incorporating both sequence and structural encoders.
Sequence encoder.
The amino acid sequence encodes the primary information of a protein and serves as the fundamental input for the UniProt-centric binding site detection. First, an input embedding module generates the initial per-residue features through: (1) learnable embeddings for the 21 amino acid types (20 standard amino acids plus an "unknown" category); (2) sinusoidal positional embedding (attention2017, ) for residue indices; (3) pre-trained ESM-2 (esm2021, ) protein embeddings. These combined features are then processed by a stack of Transformer encoder layers to capture residue-residue interactions and global sequence patterns.
Structural encoder.
Protein structure serves as the most critical input for structure-based tasks, including ligand binding site detection. Recent advances have proposed various structural feature extraction approaches, including hand-crafted algorithms (p2rank2018, ), CNN-based methods (deeppocket2022, ; deepsite2017, ) and graph neural network approaches (equipocket2023, ; vnegnn2024, ). To demonstrate the generality and effectiveness of our approach, we utilize GearNet-Edge (gearnet2023, ), a standard E(3)-invariant GNN model, without introducing any custom architecture or specialized feature engineering.
Given a protein $P$, the protein structure is represented as a residue-level relational graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{R})$, where $\mathcal{V}$ and $\mathcal{E}$ represent the set of nodes and edges respectively, and $\mathcal{R}$ is the set of edge types. Each node in the protein graph represents the alpha carbon of a residue, while sequential edges, radius edges and $K$-nearest neighbor edges are considered in the graph. Based on the defined protein graph, the node features are updated through the relational graph convolution layers (relationgraph2018, ) as follows:
$u_{i}^{(l)}=\text{ReLU}\left(\text{BN}\left(\underset{r\in\mathcal{R}}{\Sigma}W_{r}\underset{j\in\mathcal{N}_{r}(i)}{\Sigma}h_{j}^{(l-1)}\right)\right),\,h_{i}^{(l)}=h_{i}^{(l-1)}+u_{i}^{(l)}$
where $h_{i}^{(l)}$ represents the feature of node $i$ at the $l$-th layer, $\mathcal{N}_{r}(i)=\{j\in\mathcal{V}|(j,i,r)\in\mathcal{E}\}$ denotes the neighborhood of node $i$ with the edge type $r$, and $W_{r}$ is the convolutional kernel matrix shared within the edge type $r$. Specifically, BN represents a batch normalization layer and ReLU denotes the ReLU activation function.
Unlike conventional binding site detection methods that rely on structural input, our framework allows the structural encoder to be optionally included. When incorporated, the structural features are concatenated with sequence features and projected via a linear layer to match the decoder’s channels.
Transformer decoder.
The decoder comprises multiple Transformer decoder layers (attention2017, ) that simultaneously process $N$ embeddings of dimension $d_{\text{model}}$, $\mathcal{Q}\in\mathbb{R}^{N\times d_{\text{model}}}$, through multi-head self-attention and cross-attention mechanisms. Following established practices in (detr2020, ; maskformer2021, ), these input embeddings are learnable positional embeddings which we refer to as site queries. The attention mechanisms enable the decoder to perform global reasoning over all potential binding sites while incorporating contextual information from the residue-level protein features output by the encoder.
Segmentation module.
We process the $N$ site queries through a linear classifier followed by the softmax activation to generate class probabilities $\{p_{i}\in\Delta^{2}\}_{i=1}^{N}$. Here, the classifier predicts an additional $\emptyset$ (non-binding) category to indicate when a query does not correspond to any actual binding site. For mask prediction, the site queries $\mathcal{Q}\in\mathbb{R}^{N\times d_{\text{model}}}$ are converted to $N$ mask embeddings $\mathcal{E}_{\text{mask}}\in\mathbb{R}^{N\times d_{\text{model}}}$ by a Multilayer Perceptron (MLP). Finally, the binary mask prediction for each query is computed via dot-production between the $i$-th mask embedding and the residue-level protein features $\mathcal{F}\in\mathbb{R}^{L\times d_{\text{model}}}$, followed by a sigmoid activation:
$m_{i}[j]=\text{sigmoid}\left(\mathcal{E}_{\text{mask}}[i,:]\cdot\mathcal{F}[j,:]^{T}\right)$
4 Rethinking the Evaluation Metrics for Binding Site Detection
DCC (Distance between the predicted binding site center and the true binding site center) and DCA (Shortest distance between the predicted binding site center and any heavy atom of the ligand) are the two most widely-used metrics for binding site detection. A binding site prediction is considered successful when its DCC/DCA value is below a predetermined threshold. Previous works (p2rank2018, ; deepsite2017, ; deepsurf2021, ; grasp2024, ) quantify prediction performance via the Success Rate, defined as the ratio of successful predictions to the total number of ground truth sites:
[…]
However, these metrics suffer from two critical limitations: (Limitation 1) They disregard the prediction scores or ranks, and predictions may be double-counted due to the absence of proper matching criteria (Figure 4 A). (Limitation 2) They only evaluate the center of binding sites and are ligand-dependent (DCC typically considers the ligand center as site center). Since different ligands can bound to one binding site, relying solely on ligand-centered evaluation causes these metrics to completely miss key structural properties such as the shape, size, and residue composition of the binding site. It leads to evaluation failures in certain scenarios (Figure 4 B-D), disregard the crucial information required for downstream tasks (Appendix A).
Previous works have recognized these limitations. To address Limitation 1, a common approach is to calculate the DCC/DCA for either the top-$\bm{n}$ or top-($\bm{n}$+2) predicted binding sites (p2rank2018, ; equipocket2023, ; grasp2024, ), where $\bm{n}$ is the number of ground truth sites. For Limitation 2, DeepSurf (deepsurf2021, ) and Utgés et al. (cmp2024, ) propose computing the IoU between the predicted and ground truth binding site residues. Given two binding sites $m_{A},m_{B}\in\{0,1\}^{L}$, where $L$ is the protein sequence length, the IoU of two sites is defined as:
$\text{IoU}(m_{A},m_{B})=\frac{\text{sum}(m_{A}\&m_{B})}{\text{sum}(m_{A}|m_{B})},\text{where }m_{A},m_{B}\in\{0,1\}^{L}$
However, these methods fail to address the core issues, as they still lack proper matching between predicted and ground truth sites, and the top-$\bm{n}$ or top-($\bm{n}$+2) metrics introduce information leakage.
To overcome these limitations, we propose to calculate the Average Precision (AP) metric based on the residue-level IoU. We calculate AP as follows: First, we sort all predictions by confidence scores. Then, we match each ground truth site to the predicted site with the highest score and residue-level IoU above a predetermined threshold, enforcing a one-to-one assignment constraint. Finally, we compute AP as the area under the interpolated precision-recall curve following COCO evaluation protocols (coco2014, ), which is widely used in Object Detection. The AP metric offers two significant advantages: (1) the residue-level IoU enables accurate shape and size comparison between binding sites; (2) the one-to-one matching scheme inherently prevents double-counting of predictions.
Figure 4: DCC/DCA failure cases. (A) Repeated counting of the same predicted site since absence of matching. (B) Different ligands bound to the same site lead to deviations in DCC/DCA calculations. (C-D) Failed predictions classified as successful by DCC/DCA but below the IoU threshold.
5 Experiments
5.1 Settings
Dataset.
UniSite-DS is used for training and validation. We employ MMSeq2 (mmseqs2018, ) to ensure that no test UniProt sequence has similarity above 0.9 to any sequence in the training set. For each UniProt sequence, we select the PDB structure with the highest sequence identity as the representative structure. Additionally, we compare UniSite with baseline methods on widely-used binding site benchmark datasets, HOLO4K (p2rank2018, ) and COACH420 (p2rank2018, ). Following DeepSurf (deepsurf2021, ) and EquiPocket (equipocket2023, ), we use the mlig subsets of HOLO4K and COACH420 for evaluation. Since our models are trained under a UniProt-centric schema, we only consider single-chain structures, denoting the test datasets as HOLO4K-sc and COACH420 (all structures in COACH420 are originally single-chain). All test UniProt entries are strictly excluded from the training set. More details are provided in Appendix D.
Implementation details.
We set $d_{\text{model}}=256$ by default. The transformer encoder consists of 6 standard Transformer encoder layers with a feed-forward dimension 1024 and the dropout rate of 0.1. We employ 6 Transformer decoder layers following the architecture of DETR  (detr2020, ). By default, We use 32 site queries, where each query is associated with a learnable positional encoding and a zero-initialized query embedding. The multi-layer perceptron in the segmentation module consists of 2 hidden layers with 256 channels. For mask prediction, we use a combination of BCE loss and dice loss (diceloss2016, ):
$\mathcal{L}_{\text{mask}}=\lambda_{\text{bce}}\mathcal{L}_{\text{bce}}+\lambda_{\text{dice}}\mathcal{L}_{\text{dice}},$
where $\lambda_{\text{bce}}=\lambda_{\text{dice}}=5.0$. The classification loss weight $\lambda_{\text{cls}}$ is set to $2.0$, and we downweight the classification loss by a factor of 10 when $c_{i}^{gt}=\emptyset$ to mitigate class imbalance. Following DETR detr2020 , we apply segmentation modules which share the same weights after each decoder layer, and supervise their predictions by the set prediction loss. We optimize the model using the AdamW optimizer (adamw2019, ) with a learning rate of $1.0\times 10^{-4}$ and a weight decay factor of $0.05$. For the structural encoder, we implement the GearNet-Edge network following the origin paper (gearnet2023, ) without specialized feature engineering. Notably, we train the GearNet-Edge network from scratch rather than loading pre-trained weights. All models are trained on 8 NVIDIA RTX 4090 GPUs.
Our method predicts the associated residues along with a confidence score for each binding site. For applications requiring binding site centers, we compute these as the centroids of the convex hull encompassing all atoms within each predicted binding site grasp2024 ; convex1970 .
Table 1: Results on UniSite-DS. We highlight the top two performing methods for each metric in bold. ${}^{\text{a}}$ Fpocket-rescore denotes sites initially predicted by Fpocket and subsequently rescored by P2Rank. ${}^{\text{b}}$ VN-EGNN only outputs centers of predicted sites. For each center, We include the residues within a 9Å radius, which has the best AP performance (Appendix E).
Method Type Input $\text{AP}_{0.3}$ $\text{AP}_{0.5}$
Fpocket (fpocket2009, ) Geometry-based structure 0.1836 0.1017
[…]
P2Rank (p2rank2018, ) structure 0.5056 0.2157
DeepPocket (deeppocket2022, ) CNN-based structure + Fpocket result 0.4273 0.2334
GrASP (grasp2024, ) GNN-based structure 0.4469 0.2848
[…]
UniSite-1D Ours sequence 0.5121 0.3033
UniSite-3D structure 0.5603 0.3835
Evaluation metrics.
For comprehensive evaluation, we compare IoU-based AP and traditional DCC/DCA metrics in HOLO4K-sc and COACH420. Since a Uniprot-centric data entry can contain ligands from multiply PDB structures, coordinate-dependent metrics like DCC and DCA become unsuitable due to potential inconsistencies in ligand spatial arrangements. Consequently, we merely use IoU-based AP metric on UniSite-DS. Following EquiPocket (equipocket2023, ), we set the DCC/DCA threshold to 4Å, and compute the DCC/DCA success rate of top-$\bm{n}$ predictions. We calculate AP using IoU thresholds of $0.3$ and $0.5$.
5.2 Results on UniSite-DS
The results on UniSite-DS are shown in Table 1. The geometry-based method Fpocket (fpocket2009, ) exhibits inferior performance since it merely considers the geometry and electronegativity. P2Rank (p2rank2018, ) achieves better results by extracting the protein surface features with Random Forest. DeepPocket (deeppocket2022, ) utilizes 3D-CNN to rescore and refine Fpocket predictions, improving the preformance in AP0.5. Notably, Fpocket-rescore, which combines Fpocket’s initial predictions with P2Rank’s re-ranking, surpasses both P2Rank and DeepPocket, highlighting the importance of proper scoring in binding site detection as captured by our AP metric. For graph models, GrASP (grasp2024, ) achieves further improvements in AP0.5 by employing graph attention networks. However, VN-EGNN (vnegnn2024, ) performs poorly under AP metrics, because it only outputs predicted binding site centers, discarding structural properties (shape and size) or residue identification. For evaluation purposes, we include the residues within a 9Å radius of each predicted center, which has the best AP performance (Appendix E).
Trained on the UniProt-centric dataset, UniSite-1D outperforms all baseline methods without protein structure, demonstrating remarkable capability for structure-free binding site detection, particularly valuable for site-aware protein–ligand docking (Appendix A). UniSite-3D further improves the performance remarkably by incorporating structure information. The above observations not only validate the effectiveness of our methods, but also reveal the significant statistical biases inherent in previous PDB-centric datasets, which limit the performance of prior methods.
5.3 Results on HOLO4K-sc and COACH420
[…]
Table 2: Results on HOLO4K-sc and COACH420. We highlight the top two performing methods for each metric in bold. ${}^{\text{a}}$ Fpocket-rescore denotes sites initially predicted by Fpocket and subsequently rescored by P2Rank. ${}^{\text{b}}$ VN-EGNN only outputs centers of predicted sites. For each center, We include the residues within a 9Å radius, which has the best AP performance (Appendix E).
Method HOLO4K-sc COACH420
$\text{AP}_{0.3}$ $\text{DCC}_{\text{top-}n}$ $\text{DCA}_{\text{top-}n}$ $\text{AP}_{0.3}$ $\text{DCC}_{\text{top-}n}$ $\text{DCA}_{\text{top-}n}$
Fpocket (fpocket2009, ) 0.2711 0.3076 0.4382 0.2106 0.2708 0.4107
[…]
P2Rank (p2rank2018, ) 0.6011 0.5300 0.8188 0.6188 0.4643 0.7411
DeepPocket (deeppocket2022, ) 0.5415 0.4925 0.7369 0.5184 0.3958 0.6756
GrASP (grasp2024, ) 0.6668 0.5131 0.7416 0.7150 0.4851 0.7620
[…]
UniSite-1D (ours) 0.6867 0.5538 0.7692 0.5921 0.4554 0.7351
UniSite-3D (ours) 0.7091 0.5716 0.7879 0.7196 0.4702 0.7381
Table 3: Effect of sequence similarity. The second column indicates the sequence identity between training sets and test sets.
Method Similarity $\text{AP}_{0.3}$ $\text{AP}_{0.5}$
UniSite-1D <0.9 0.5121 0.3033
UniSite-3D 0.5603 0.3835
UniSite-1D <0.7 0.5056 0.2945
UniSite-3D 0.5579 0.3734
UniSite-1D <0.5 0.4338 0.2243
UniSite-3D 0.4677 0.2801
Table 4: Effect of site queries. This table shows results of UniSite-3D trained with a varying number of site queries.
# of queries $\text{AP}_{0.3}$ $\text{AP}_{0.5}$
16 0.5515 0.3795
32 0.5603 0.3835
48 0.5562 0.3861
64 0.5615 0.3867
5.4 Ablation study
Sequence similarity.
Both the structure and function of proteins diverge with the decreasing of sequence similarity. It is necessary to evaluate our protein ligand binding site detection method across varying similarity thresholds. We employ MMSeqs2 (mmseqs2018, ) to partition UnSite-DS with three similarity thresholds: 0.5, 0.7 and 0.9, ensuring that no test protein exceeds the corresponding similarity to any training protein. Compared to threshold 0.9, UnSite-1D/3D exhibit only a slight decrease under threshold 0.7 (Table 4), indicating that our methods possess generalization ability. As proteins with sequence similarity below 0.5 typically belong to evolutionarily distant families and exhibit markedly different structural folds, we observe significant AP performance degradation for UniSite-1D and UniSite-3D under threshold 0.5.
Number of site queries.
As shown in Table 4, UniSite-3D exhibits stable performance across varying numbers of site queries. we select 32 queries as our default configuration, considering both the computation cost and the coverage of ground truth binding sites (99.5% proteins in UniSite-DS have sites less than 20).
6 Conclusion
In this paper, we introduce UniSite-DS, the first UniProt-centric dataset of protein ligand binding sites, which systematically integrates all ligand binding sites across multiple PDB structures for each unique protein. UniSite-DS corrects the statistical bias in previously available PDB-centric datasets and methods, while significantly broadening the available data. To amend the discontinuous workflows in existing binding site detection methods, we proposed UniSite-1D/3D, two end-to-end methods supervised by set prediction loss with bijective matching. In addition, we introduce IoU-based AP as a more accurate evaluation metric. Extensive experiments on UniSite-DS and several benchmark datasets demonstrate that our frameworks achieve superior performance, and the IoU-based AP metric can provide a more accurate reflection of binding site prediction quality.
Appendix A Impact of Binding Site Accuracy on Downstream Tasks
The accuracy of protein–ligand binding site detection is critical for downstream tasks. Taking molecular docking as an example (Figure S1), a comparison of different methods leads to the following conclusions: (1) The definition of binding site residues can strongly impact the docking performance, highlighting the importance of our proposed IoU-based AP metric. (2) Docking methods that do not specify binding sites (blind docking) show significant performance improvements when provided with binding site information, emphasizing the importance of accurate binding site identification.
Figure S1: The significant impact of binding site detection on molecular docking. (A) Gold [47] defines the binding site using a sphere. (B) AutoDock Vina [48] defines the binding site using a cube. (C) DeepDock [49] and (D) Uni-Mol [50] identify the binding site by applying a fixed radius around the ligand. (E) Docking success rates on the PoseBusters dataset under different binding site configurations. Docking success rate is defined as the proportion of predictions with an RMSD less than 2Å. Data sourced from  [51, 52, 53, 54].
Appendix B Case Study of Classical Methods
Figure S2: Case study of classical methods. For UniProt ID Q7YYQ9, the two ground truth binding sites are represented by dark blue and orange meshes, respectively. All predicted binding sites are shown as surfaces. (A) The two ground truth binding sites are colored in dark blue and orange. (B) The two binding sites predicted by our UniSite-3D method are colored in cyan and purple. (C) The single binding site predicted by the GNN-based method GrASP [14] is colored in green. (D) The single binding site predicted by the classical machine learning method P2Rank [11] is colored in red. (E) The single binding site predicted by the geometry-based method Fpocket [8] is colored in cyan.
As discussed in Section 1, 2, previous approaches are developed based on PDB-centric datasets, which introduce substantial statistical biases. We conduct a case study of these methods, and the results (Figure S2) indicate that, for multi-site proteins, existing approaches are weak in distinguishing between different ligand binding sites. This limitation motivated us to investigate the problems in existing approaches and to develop a new methodology.
Appendix C Related Work
Over the past several decades, numerous methods have been developed for detecting protein–ligand binding sites, accompanied by advances in techniques leveraging the geometric, physical and chemical features of proteins.
Early methods relied on traditional computational algorithms. Since most binding sites show up as cavities in protein 3D structures, geometry-based methods (Fpocket [8], LigSite [55]) identify and rank these hollow cavities through hand-crafted features like alpha spheres [56]. Template-based methods (FINDSITE [9] and LIBRA [57]) predict ligand binding sites by comparing the query protein with templates from known protein structure database. These methods typically generate a large number of predicted sites while performing poorly in ranking them.
Subsequent approaches like PRANK [58] and P2Rank [11] employ traditional machine learning methods, particularly Random Forest. Based on the predictions of Fpocket [8], PRANK assigns "ligandibility" scores, which denotes ligand binding potential, to candidate sites. P2Rank is a widely used method which integrates the geometric features of the protein surface with Random Forest Algorithm.
In recent years, deep learning methods have emerged for protein–ligand binding site detection. CNN-based approaches [12, 36, 40] treat protein structures as 3D images, applying 3D convolutional neural networks similar to those used in computer vision. Alternatively, GNN-based methods [15, 14, 13] utilize graph neural networks by constructing graphs incorporating both geometric and chemical features of proteins. Despite their improved performance, these methods typically adopt discontinuous workflows: they first perform semantic segmentation to generate binary masks of potential binding residues/atoms, then cluster these masks into discrete binding sites. This fragmented pipeline heavily depends on post-processing (e.g., clustering algorithms), inherently limiting end-to-end optimization and struggle with overlapping binding sites.
Appendix D External Datasets
scPDB [16] is a famous dataset for protein–ligand binding site detection, commonly employed for training and validation in recent studies ( [15, 14, 13]). scPDB provides both protein and ligand structures, accompanied by the structures of binding site extracted via VolSite [59]. Notably, only one binding site and one corresponding ligand are annotated for each data entry. In this work, we use the 2017 release of scPDB, which contains 17,594 structures and 5,550 unique proteins. (Source: http://bioinfo-pharma.u-strasbg.fr/scPDB/)
PDBBind [17] is a widely used dataset to study protein–ligand interaction, especially for protein–ligand docking [51, 60]. Similar to scPDB, PDBBind annotates one ligand structure and one binding site structure in each data entry. In this paper, we use the general set of v2020, the latest academic-free edition, which comprises 19,443 structures and 3,888 unique proteins. (Source: http://www.pdbbind.org.cn/download/)
HOLO4K and COACH420 are two benchmark datasets utilized for protein–ligand binding site detection. Follow VN-EGNN [15], EquiPocket [13] and GrASP [14], we employ the mlig subsets of these two dataset, which contain explicitly specified relevant ligands. HOLO4K-mlig comprises 3,204 structures and 1,259 unique proteins, while COACH420-mlig covers 284 structures and 265 unique proteins. (Source: https://github.com/rdk/p2rank-datasets)
Appendix E The AP Evaluation of VN-EGNN
Since VN-EGNN [15] outputs only the centers of predicted binding sites, it is non-trivial to identify the binding residues. In order to evaluate the IoU-based AP, we include the residues with a fixed radius for each predicted center (Figure S3). As shown in Table S1 , the radius of 9Å has the best performance across all datasets.
Figure S3: Binding sites derived from the predicted center by VN-EGNN using different radii. Binding site residues are visualized using surface representation: (A) residues within a 4.5Å radius; (B) residues within a 6Å radius; (C) residues within a 9Å radius; (D) residues within a 12Å radius.
Table S1: AP evaluation results about VN-EGNN with different radius.
Radius UniSite-DS HOLO4K-sc COACH420
$\text{AP}_{0.3}$ $\text{AP}_{0.5}$ $\text{AP}_{0.3}$ $\text{AP}_{0.5}$ $\text{AP}_{0.3}$ $\text{AP}_{0.5}$
4.5Å 0.0054 0.0004 0.0049 0.0001 0.0072 0.0007
0.0894 0.0088 0.1290 0.0172 0.1814 0.0189
0.1621 0.0705 0.2606 0.1346 0.2637 0.1138
12Å 0.0087 0.0010 0.1411 0.0014 0.1444 0.0013
Appendix F Full Results on HOLO4K-sc and COACH420
Table S2: Full results on HOLO4K-sc. We highlight the top two performing methods for each metric in bold. ${}^{\text{a}}$ Fpocket sites rescored by P2Rank. ${}^{\text{b}}$ Residues within 9Å of each VN-EGNN predicted center, which has the best AP performance (Appendix E).
Method HOLO4K-sc
$\text{AP}_{0.3}$ $\text{AP}_{0.5}$ $\text{DCC}_{\text{top-}n}$ $\text{DCC}_{\text{top-}n+2}$ $\text{DCA}_{\text{top-}n}$ $\text{DCA}_{\text{top-}n+2}$
Fpocket [8] 0.2711 0.1488 0.3076 0.4181 0.4382 0.5941
[…]
P2Rank [11] 0.6011 0.2625 0.5300 0.5623 0.8188 0.8652
DeepPocket [12] 0.5415 0.2891 0.4925 0.5478 0.7369 0.7851
GrASP [14] 0.6668 0.4126 0.5131 0.5267 0.7416 0.7612
[…]
UniSite-1D (ours) 0.6867 0.4595 0.5538 0.6400 0.7692 0.8305
UniSite-3D (ours) 0.7091 0.5446 0.5716 0.6470 0.7879 0.8422
Table S3: Full results on COACH420. We highlight the top two performing methods for each metric in bold. ${}^{\text{a}}$ Fpocket sites rescored by P2Rank. ${}^{\text{b}}$ Residues within 9Å of each VN-EGNN predicted center, which has the best AP performance (Appendix E).
Method COACH420
$\text{AP}_{0.3}$ $\text{AP}_{0.5}$ $\text{DCC}_{\text{top-}n}$ $\text{DCC}_{\text{top-}n+2}$ $\text{DCA}_{\text{top-}n}$ $\text{DCA}_{\text{top-}n+2}$
Fpocket [8] 0.2106 0.1219 0.2708 0.3750 0.4107 0.5714
[…]
P2Rank [11] 0.6188 0.2618 0.4643 0.5000 0.7411 0.8034
DeepPocket [12] 0.5184 0.2512 0.3958 0.4821 0.6756 0.7560
GrASP [14] 0.7150 0.4914 0.4851 0.4970 0.7620 0.7917
[…]
UniSite-1D (ours) 0.5921 0.2998 0.4554 0.5238 0.7351 0.8006
UniSite-3D (ours) 0.7196 0.3977 0.4702 0.5387 0.7381 0.8095